A dress was first marked down 20% off original price. Then marked down 25
% more off the first sale price and finally 30% off the second sale price. It was sold for 63$ What was the original price? There is no textbook the teacher made it up.

Answer by checkley71(8403)

GIVEN THE ORIGINAL PRICE WAS X THE X*.8*.75*.7=63 OR .42X=63 OR X=63/.42 OR
X=150
PROOF 150-(.2*150)=150-30=120 THEN 120-(.25*120)=120-30=90 OR
90-(.3*90)=90-27=63